# HELM: Hierarchical Encoding for mRNA Language Modeling

### v1

@@ Abstract
Messenger RNA (mRNA) plays a crucial role in protein synthesis, with its codon structure directly impacting biological properties. While Language Models (LMs) have shown promise in analyzing biological sequences, existing approaches fail to account for the hierarchical nature of mRNA’s codon structure. We introduce Hierarchical Encoding for mRNA Language Modeling (HELM), a novel pre-training strategy that incorporates codon-level hierarchical structure into language model training. HELM modulates the loss function based on codon synonymity, aligning the model’s learning process with the biological reality of mRNA sequences. We evaluate HELM on diverse mRNA datasets and tasks, demonstrating that HELM outperforms standard language model pre-training as well as existing foundation model baselines on six diverse downstream property prediction tasks and an antibody region annotation tasks on average by around 8%. Additionally, HELM enhances the generative capabilities of language model, producing diverse mRNA sequences that better align with the underlying true data distribution compared to non-hierarchical baselines.
[…]
@@ 1 Introduction
RNA analysis is becoming increasingly important in molecular biology (Liu et al., 2023; Fu, 2014). Messenger RNA (mRNA) is of particular interest due to its unique role in protein synthesis  (Sahin et al., 2014). mRNA consists of triplets of nucleotides, called codons, which directly translate to protein amino acids through a surjective many-to-one mapping (Figure 1 left). This results in multiple synonymous mRNAs that encode identical amino acid sequences while exhibiting distinct physical and biological properties at the nucleotide level (Buhr et al., 2016). The role of mRNA, serving as an intermediary between DNA and proteins, underlies its therapeutic potential in applications such as vaccines and gene therapies (Caplen, 2004; Pardi et al., 2018), while also presenting challenges for analysis and optimization (Liu et al., 2021).
@@
To demonstrate the practical advantages of HELM, we focus on the important domain of antibody mRNAs where codon distribution can significantly impact physical and therapeutic properties  (Zhang et al., 2023). We conduct comprehensive experiments, evaluating our hierarchical model against its non-hierarchical counterparts pre-trained with natural-language CLM and MLM as well as against other state-of-the-art RNA foundation models. Our results show that hierarchical pre-training yields significantly better representations of antibody-encoding mRNA and even excels on non-antibody mRNA data. Our hierarchy-aware model outperforms non-hierarchical counterparts pre-trained with natural-language CLM and MLM as well as other state-of-the-art RNA foundation models by around 8% on average across six diverse downstream property prediction tasks and an antibody region annotation tasks while requiring up to 2 times fewer model parameters. Beyond its excellent discriminative performance, we demonstrate that HELM-based CLM pre-training yields strong generative models capable of producing diverse mRNA sequences while preserving closer alignment with the true data distribution. Ultimately, HELM addresses the limitations of current bio-LMs by effectively capturing the hierarchical structure of mRNA as a strong biological prior, leading to improved performance in both predictive and generative tasks while requiring minimal modification of the pre-training.
@@
  • We provide a consistent comparison of various tokenization methods, hierarchical and standard pre-training strategies, and different LM architectures to facilitate further development of mRNA models.
@@ 2 Related works
Recent advancements in LMs have significantly improved the analysis of biological sequences such as DNA, RNA, and proteins (Laurent et al., 2024). The standard approach for bio-language modeling adopts pre-training strategies from natural language; state-of-the-art protein or DNA LMs rely on Masked Language Modeling (Liu et al., 2020c; Lin et al., 2023; Dalla-Torre et al., 2023) or Causal Language Modeling (Brown, 2020; Lin et al., 2023) and pre-train representations on large bio-databases (Bairoch et al., 2005; Suzek et al., 2007; Consortium et al., 2013; Schoch et al., 2020). In addition, the modeling choices of pre-training bio-LMs involve network architecture and tokenization strategy. Regarding architectures, transformer-based models have been widely adopted (Elnaggar et al., 2021; Lin et al., 2023), while recent works explore structured state-space (Peng et al., 2024; Sgarbossa et al., 2024) and long-convolutional models (Nguyen et al., 2024b; a; Moskalev et al., 2024) to better handle long-range dependencies in biological sequences. Tokenization strategies vary from character-level approaches (Elnaggar et al., 2021; Lin et al., 2023) to k-mer tokenization (Dalla-Torre et al., 2023) and Byte Pair Encoding  (Zhou et al., 2023). These existing pre-training and modeling choices primarily adapt natural language paradigms to biological sequences, overlooking the incorporation of inherent biological structures crucial for mRNA analysis. Moreover, the lack of consistent comparisons between these various approaches hinders the development of effective mRNA LMs. Our work addresses these limitations with HELM, a pre-training strategy incorporating mRNA’s hierarchy into the training, and we also provide analysis of various pre-training and modeling choices to facilitate further development of mRNA LMs.
@@ 3 Hierarchical Cross-Entropy Loss
We adapt the Hierarchical Cross-Entropy to incorporate the biological structure of mRNA into Masked Language Modeling and Causal Language Modeling pre-training strategies. For the hierarchical MLM, we randomly mask a portion of codons in the input sequence and train a model to predict the masked codons supervising it by the HXE loss. In the hierarchical CLM, a model predicts the next codon given the previous ones with the HXE loss function.
The hierarchical structure of mRNA informs the HXE objective, allowing the model’s output token probabilities to be factorized according to the tree structure. Then, the HXE loss uses Eq.1 and Eq.2 weighting the MLM or CLM errors differently based on the codon’s position in the hierarchy. We employ a practical weighting function $\lambda(C)=\exp(-\alpha h(C))$ where $h(C)$ is the height of the node $C$ in the hierarchy and ($\alpha>0$). The value of $\alpha$ determines how much weight is given to information at different hierarchical levels. This weighting function is applied in both MLM and CLM pre-training, allowing HELM to capture both local codon-level information and the global hierarchical structure of the genetic code. By incorporating this hierarchical loss into pre-training, HELM encourages the model to learn biologically meaningful representations of mRNA sequences, which closely aligns the model’s learning with the biological structure of mRNA sequences. Note that our approach does not require modifying the architecture or tokenization of a model and permits operating on the standard codon vocabulary. mRNA hierarchy modeling with HXE will be referred as HELM from here on.
@@ 4 Experiments
For downstream evaluation, we use the following six mRNA datasets, including antibody (Ab) encoding mRNA and general mRNA sequences:
  • Ab1 (Anonymous, 2024) includes 1200 Ab-mRNA sequences with protein expression labels.
  • Ab2 (Anonymous, 2024) contains 3442 Ab-mRNA sequences with protein expression labels.
@@
We ran all models for $3$ random data splits (train:val:test split of 70:15:15) and we report average performance across splits. When evaluating property prediction tasks, we use Spearman rank correlation as a standard metric under commonly used probing methodology (Marquet et al., 2022; Chen et al., 2024; Outeiral & Deane, 2024; Harmalkar et al., 2023) with a TextCNN (Kim, 2014) head to assess representation quality and transferability. For generative evaluation, we take the metrics used to evaluate generative models for DNA (Stark et al., 2024; Shuai et al., 2023) and we adopt them for mRNA data. For sequence region annotation, we report region annotation accuracy. We provide comprehensive experimental details in Appendix A.3.
@@ 4.1.1 Experimental details
We evaluate various state-of-the-art architectures for sequence modeling: Transformer (Radford et al., 2019), Hyena (Poli et al., 2023), and Mamba (Gu & Dao, 2023). Transformers support both MLM and CLM objectives, while Hyena and Mamba are primary designed for CLM. All models use 50M parameters, balancing performance and efficiency. We found that models of this scale can outperforms larger existing models while maintaining reasonable run-times (see Appendix A.7 and Table 8). Detailed pre-training information is available in Appendix A.2.
@@
Inspired by protein and DNA language modeling, we explored various tokenization strategies for mRNA sequences: nucleotide, codon-level, and 6-mer. We observed that codon-level tokenization consistently outperforms other strategies (results in Appendix Sec. A.4). We hypothesize that the superior performance of codon tokenization stems from codons’ role as the fundamental units of the genetic code, directly mapping to amino acids during protein synthesis. In addition, the codon level tokenization allows to naturally capture the variability in wobbling nucleotides, where changes in the third nucleotide do not always alter the amino acid produced (Barricelli, 1977; Outeiral & Deane, 2024). Here on, we adopt codon-level tokenization as the default strategy for our models. This way, our vocabulary includes $644^{3}=$ 64 codon tokens, representing all possible nucleotide combinations, plus special tokens for start, stop, padding, masking, and unknown characters.
@@ 4.1.2 Results
We first compare the performance of the transformer-based mRNA language models to state-space Mamba and long-convolutional Hyena-based architectures. In Table 1, we can see that transformer-based models outperform Hyena and Mamba on 4 out of 6 datasets, with the transformer CLM trailing closely behind Mamba on the remaining 2 datasets. The superior performance of transformers is likely due to the relatively short sequence lengths ($\sim$ 1000 tokens) of mRNA sequences in both pre-training and downstream datasets. Transformers, with their explicit global attention mechanism between each pair of tokens, excel at capturing fine-grained interactions in these relatively short sequences. In contrast, Hyena and Mamba architectures are optimized for handling much longer sequences via efficient convolutional or state-space mechanisms and may not fully leverage their strengths in this shorter sequence regime owing to their implicit handling of sequence dependencies, despite performing competitively. Given that transformer-based models generally show better performance than Hyena and Mamba models, we select transformer-based MLM and CLM models for all subsequent experiments.
@@
Model Ab1 Ab2 MLOS iCodon Tc-Riboswitches mRFP
@@
Transformer XE (MLM) 0.748 0.599 0.653 0.503 0.569 0.753
Transformer XE (CLM) 0.752 0.597 0.611 0.498 0.531 0.815
Hyena XE 0.743 0.594 0.623 0.526 0.517 0.844
Mamba XE 0.742 0.585 0.620 0.509 0.520 0.823
@@ 4.2 HELM improves mRNA property prediction
Next, we evaluate the effect of learning hierarchical mRNA encoding with HELM for the mRNA property prediction tasks. As can be observed from Table 2, training with biological hierarchical prior consistently improves model performance for both MLM and CLM objectives on all six datasets compared to baseline models trained with standard non-hierarchical cross-entropy. For causal language modeling, HELM-CLM outperforms the non-hierarchical XE-CLM model in five out of six datasets. For the masked language modeling, HELM-MLM outperforms the non-hierarchical XE-MLM model in all datasets. Within hierarchy-aware models, we again observe the same trend that MLM and CLM HELM models perform comparably for most tasks with each performing best on three of the six datasets.
@@
Model Ab1 Ab2 MLOS iCodon Tc-Riboswitches mRFP
Transformer XE (MLM) 0.748 0.599 0.653 0.503 0.569 0.753
Transformer HELM (MLM) 0.767 0.609 0.701 0.525 0.626 0.822
Transformer XE (CLM) 0.752 0.597 0.611 0.498 0.531 0.815
Transformer HELM (CLM) 0.760 0.614 0.592 0.529 0.619 0.849
@@
To evaluate synonymous codon usage bias for any dataset, we first calculate the distribution of synonymous codons for each amino acid and then compute the entropy of codon usage for the five most frequent amino acids. Lower entropy indicates less uniform codon distribution hence the preference towards specific codons or codon bias. Fig. 3 and Appendix Table 7 demonstrate that MLOS, Tc-Riboswitches, and mRFP datasets have lower entropy, suggesting stronger codon bias, which aligns with the larger improvements seen in these datasets using HELM models. These findings suggest that HELM’s hierarchical approach is particularly beneficial for datasets with pronounced codon usage biases, highlighting its potential for improved mRNA modeling in biologically diverse contexts.
@@
Evaluating hierarchical encoding through synonymous sequence clustering
MLM Silh. ($\uparrow$) CLM Silh. ($\uparrow$)
XE 0.74 0.68
HELM 0.91 0.83
MLM Acc. ($\uparrow$) CLM Acc. ($\uparrow$)
XE 78.68 67.51
HELM 83.39 73.39

### v2

@@ Abstract
Messenger RNA (mRNA) plays a crucial role in protein synthesis, with its codon structure directly impacting biological properties. While Language Models (LMs) have shown promise in analyzing biological sequences, existing approaches fail to account for the hierarchical nature of mRNA’s codon structure. We introduce Hierarchical Encoding for mRNA Language Modeling (HELM), a novel pre-training strategy that incorporates codon-level hierarchical structure into language model training. HELM modulates the loss function based on codon synonymity, aligning the model’s learning process with the biological reality of mRNA sequences. We evaluate HELM on diverse mRNA datasets and tasks, demonstrating that HELM outperforms standard language model pre-training as well as existing foundation model baselines on seven diverse downstream property prediction tasks and an antibody region annotation tasks on average by around 8%. Additionally, HELM enhances the generative capabilities of language model, producing diverse mRNA sequences that better align with the underlying true data distribution compared to non-hierarchical baselines.
@@ 1 Introduction
RNA analysis is becoming increasingly important in molecular biology (Liu et al., 2023; Fu, 2014). Messenger RNA (mRNA) is of particular interest due to its unique role in protein synthesis  (Sahin et al., 2014). mRNA consists of triplets of nucleotides, called codons, which directly translate to protein amino acids through a surjective many-to-one mapping (Figure 1 left). This results in multiple synonymous mRNAs that encode identical amino acid sequences while exhibiting distinct physical and biological properties at the nucleotide level (Buhr et al., 2016). The role of mRNA, serving as an intermediary between DNA and proteins, underlies its therapeutic potential in applications such as vaccines and gene therapies (Caplen, 2004; Pardi et al., 2018), while also presenting challenges for analysis and optimization (Liu et al., 2021; Xu et al., 2025).
@@
To demonstrate the practical advantages of HELM, we focus on the important domain of antibody mRNAs where codon distribution can significantly impact physical and therapeutic properties  (Zhang et al., 2023). We conduct comprehensive experiments, evaluating our hierarchical model against its non-hierarchical counterparts pre-trained with natural-language CLM and MLM as well as against other state-of-the-art RNA foundation models. Our results show that hierarchical pre-training yields significantly better representations of antibody-encoding mRNA and even excels on non-antibody mRNA data. Our hierarchy-aware model outperforms non-hierarchical counterparts pre-trained with natural-language CLM and MLM as well as other state-of-the-art RNA foundation models by around 8% on average across seven diverse downstream property prediction tasks and an antibody region annotation tasks while requiring up to 2 times fewer model parameters. Beyond its excellent discriminative performance, we demonstrate that HELM-based CLM pre-training yields strong generative models capable of producing diverse mRNA sequences while preserving closer alignment with the true data distribution. Ultimately, HELM addresses the limitations of current bio-LMs by effectively capturing the hierarchical structure of mRNA as a strong biological prior, leading to improved performance in both predictive and generative tasks while requiring minimal modification of the pre-training.
@@
  • We provide a consistent comparison of various tokenization methods, hierarchical and standard pre-training strategies, and different LM architectures to facilitate further development of mRNA models. We also introduce a large high-quality curated mRNA dataset to help further research for mRNA language modeling.
@@ 2 Related works
Recent advancements in LMs have significantly improved the analysis of biological sequences such as DNA, RNA, and proteins (Laurent et al., 2024; Prakash et al., 2024). The standard approach for bio-language modeling adopts pre-training strategies from natural language; state-of-the-art protein or DNA LMs rely on Masked Language Modeling (Liu et al., 2020c; Lin et al., 2023; Dalla-Torre et al., 2023) or Causal Language Modeling (Brown, 2020; Lin et al., 2023) and pre-train representations on large bio-databases (Bairoch et al., 2005; Suzek et al., 2007; Consortium et al., 2013; Schoch et al., 2020). In addition, the modeling choices of pre-training bio-LMs involve network architecture and tokenization strategy. Regarding architectures, transformer-based models have been widely adopted (Elnaggar et al., 2021; Lin et al., 2023), while recent works explore structured state-space (Peng et al., 2024; Sgarbossa et al., 2024) and long-convolutional models (Nguyen et al., 2024b; a; Moskalev et al., 2024) to better handle long-range dependencies in biological sequences. Tokenization strategies vary from character-level approaches (Elnaggar et al., 2021; Lin et al., 2023) to k-mer tokenization (Dalla-Torre et al., 2023) and Byte Pair Encoding  (Zhou et al., 2023). These existing pre-training and modeling choices primarily adapt natural language paradigms to biological sequences, overlooking the incorporation of inherent biological structures crucial for mRNA analysis. Moreover, the lack of consistent comparisons between these various approaches hinders the development of effective mRNA LMs. Our work addresses these limitations with HELM, a pre-training strategy incorporating mRNA’s hierarchy into the training, and we also provide analysis of various pre-training and modeling choices to facilitate further development of mRNA LMs.
@@ 3 Hierarchical Cross-Entropy Loss
We adapt the Hierarchical Cross-Entropy to incorporate the biological structure of mRNA into Masked Language Modeling and Causal Language Modeling pre-training strategies. For the hierarchical MLM, we randomly mask a portion of codons in the input sequence and train a model to predict the masked codons supervising it by the HXE loss. In the hierarchical CLM, a model predicts the next codon given the previous ones with the HXE loss function. The hierarchical structure of mRNA informs the HXE objective, allowing the model’s output token probabilities to be factorized according to the tree structure. Then, the HXE loss uses Eq.1 and Eq.2 weighting the MLM or CLM errors differently based on the codon’s position in the hierarchy. We employ a practical weighting function $\lambda(C)=\exp(-\alpha h(C))$ where $h(C)$ is the height of the node $C$ in the hierarchy and ($\alpha>0$). The value of $\alpha$ determines how much weight is given to information at different hierarchical levels. This weighting function is applied in both MLM and CLM pre-training, allowing HELM to capture both local codon-level information and the global hierarchical structure of the genetic code. By incorporating this hierarchical loss into pre-training, HELM encourages the model to learn biologically meaningful representations of mRNA sequences, which closely aligns the model’s learning with the biological structure of mRNA sequences. Note that our approach does not require modifying the architecture or tokenization of a model and permits operating on the standard codon vocabulary. mRNA hierarchy modeling with HXE will be referred as HELM from here on.
@@ 4 Experiments
For downstream evaluation, we use the following seven mRNA datasets, including antibody (Ab) encoding mRNA and general mRNA sequences:
  • Ab1 (Anonymous, 2024) includes 1,200 Ab-mRNA sequences with protein expression labels, which are critical for optimizing mRNA design to ensure efficient protein translation, a key factor in the development and manufacturing of mRNA vaccines, including cancer immunotherapies.
  • Ab2 (Anonymous, 2024) contains 3442 Ab-mRNA sequences with protein expression labels, acquired using a different experimental platform than Ab1.
@@
  • COVID-19 Vaccine (Wayment-Steele et al., 2022) has 2400 mRNA sequences with degradation labels relevant for mRNA vaccine formulations.
@@
For iCodon, Tc-Riboswitches, mRFP and COVID-19 Vaccine datasets, we use pre-defined splits from prior publications to ensure fair comparison. For other datasets, we apply clustering-based train/validation/test splitting (LinClust (Steinegger & Söding, 2018) similarity threshold 0.9) to prevent data leakage. We use a train/validation/test split ratio of 70:15:15. When evaluating property prediction tasks, we use Spearman rank correlation as a standard metric under commonly used probing methodology (Marquet et al., 2022; Chen et al., 2024; Outeiral & Deane, 2024; Harmalkar et al., 2023) with a TextCNN (Kim, 2014) head to assess representation quality and transferability. For generative evaluation, we take the metrics used to evaluate generative models for DNA (Stark et al., 2024; Shuai et al., 2023) and we adopt them for mRNA data. For sequence region annotation, we report region annotation accuracy. We provide comprehensive experimental details in Appendix A.3.
@@ 4.1.1 Experimental details
We evaluate various state-of-the-art architectures for sequence modeling: Transformer (Radford et al., 2019), Hyena (Poli et al., 2023), and Mamba (Gu & Dao, 2023). Transformers support both MLM and CLM objectives, while Hyena and Mamba are primary designed for CLM. All models use 50M parameters, balancing performance and efficiency. We found that models of this scale can outperforms larger existing models while maintaining reasonable run-times (see Appendix A.8). Detailed pre-training information is available in Appendix A.2.
@@
Inspired by protein and DNA language modeling, we explored various tokenization strategies for mRNA sequences: nucleotide, codon-level, and 6-mer. We observed that codon-level tokenization consistently outperforms other strategies (results in Appendix Sec. A.4). We hypothesize that the superior performance of codon tokenization stems from codons’ role as the fundamental units of the genetic code, directly mapping to amino acids during protein synthesis. In addition, the codon level tokenization allows to naturally capture the variability in wobbling nucleotides, where changes in the third nucleotide do not always alter the amino acid produced (Barricelli, 1977; Outeiral & Deane, 2024). Here on, we adopt codon-level tokenization as the default strategy for our models. This way, our vocabulary includes $644^{3}=$ 64 codon tokens for all possible nucleotide combinations, plus special tokens for start, stop, padding, masking, and unknown characters.
@@ 4.1.2 Results
We first compare the performance of the transformer-based mRNA language models to state-space Mamba and long-convolutional Hyena-based architectures. In Table 1, we can see that transformer-based models outperform Hyena and Mamba on 5 out of 7 datasets, with the transformer CLM trailing closely behind Mamba on the remaining 2 datasets. The superior performance of transformers is likely due to the relatively short sequence lengths ($\sim$ 1000 tokens) of mRNA sequences in both pre-training and downstream datasets. Transformers, with their explicit global attention mechanism between each pair of tokens, excel at capturing fine-grained interactions in these relatively short sequences. In contrast, Hyena and Mamba architectures are optimized for handling much longer sequences via efficient convolutional or state-space mechanisms and may not fully leverage their strengths in this shorter sequence regime owing to their implicit handling of sequence dependencies, despite performing competitively. Given that transformer-based models generally show better performance than Hyena and Mamba models, we select transformer-based MLM and CLM models for all subsequent experiments.
@@
Model Ab1 Ab2 MLOS iCodon Tc-Riboswitches mRFP COV-19 Vaccine
@@
Transformer XE (MLM) 0.748 0.599 0.653 0.503 0.569 0.753 0.801
Transformer XE (CLM) 0.752 0.597 0.611 0.498 0.531 0.815 0.787
Hyena XE 0.743 0.594 0.623 0.526 0.517 0.844 0.797
Mamba XE 0.742 0.585 0.620 0.509 0.520 0.823 0.741
@@ 4.2 HELM improves mRNA property prediction
Next, we evaluate the effect of learning hierarchical mRNA encoding with HELM for the mRNA property prediction tasks. As can be observed from Table 2, training with biological hierarchical prior consistently improves model performance for both MLM and CLM objectives on all six datasets compared to baseline models (denoted as XE) trained with standard non-hierarchical cross-entropy. Note that both XE and HELM have the same architecture, tokenization and pre-training dataset, thus the superior performance of HELM serves as an ablation showcasing the benefit of learning hierarchy. For causal language modeling, HELM-CLM outperforms the non-hierarchical XE-CLM model in five out of six datasets. For the masked language modeling, HELM-MLM outperforms the non-hierarchical XE-MLM model in all datasets. Within hierarchy-aware models, we again observe the same trend that MLM and CLM HELM models perform comparably for most tasks with each performing best on three of the six datasets. Results in terms of Pearson correlation and $R^{2}$ are reported in Appendix A.11 and show similar trends. Additional analysis in Appendix Sec. A.11, conducted on real-world iCodon thermostability dataset, shows that HELM also remains more robust and superior to XE, for extreme sequence lengths and high GC content. Scaling experiments are also presented in Appendix Sec. A.8 demonstrating that hierarchical encoding remains critical even at larger model sizes.
@@
Model Ab1 Ab2 MLOS iCodon Tc-Riboswitches mRFP COV-19 Vaccine
Transformer XE (MLM) 0.748 0.599 0.653 0.503 0.569 0.753 0.801
Transformer HELM (MLM) 0.767 0.609 0.701 0.525 0.626 0.822 0.833
Transformer XE (CLM) 0.752 0.597 0.611 0.498 0.531 0.815 0.787
Transformer HELM (CLM) 0.760 0.614 0.592 0.529 0.619 0.849 0.789
@@
To evaluate synonymous codon usage bias for any dataset, we first calculate the distribution of synonymous codons for each amino acid and then compute the entropy of codon usage for the five most frequent amino acids. Lower entropy indicates less uniform codon distribution hence the preference towards specific codons or codon bias. Fig. 3 and Appendix Table 5 demonstrate that MLOS, Tc-Riboswitches, and mRFP datasets have lower entropy, suggesting stronger codon bias, which aligns with the larger improvements seen in these datasets using HELM models.
In addition to entropy-based metric, we extend our analysis of codon usage bias by looking at Codon Pair Bias (CPB), which assesses the over- or under-representation of codon pairs or higher-order k-mers within a sequence compared to a background distribution (Coleman et al., 2008). Each codon k-mer is assigned a weight based on the log-ratio of its observed to expected frequency. The overall CPB score for a sequence is the mean of these weights, with positive scores indicating over-represented k-mers and negative scores reflecting under-represented ones. Fig. 3 shows average CPB scores for each dataset, reinforcing findings from the entropic analysis and reinforcing the hypothesis that HELM’s hierarchical approach is particularly beneficial for datasets with pronounced codon usage biases, highlighting its potential for improved mRNA modeling.
@@
Figure 3: Relationship between Codon Pair Bias (CPB) and % improvement over XE. Datasets with more negative CPB values (indicating stronger codon usage bias) tend to exhibit greater improvements with HELM.